# Comparing cosmic shear nulling methods for Stage-IV surveys

### v1

@@ 5 Impact of nulling strategies on the cosmic shear power spectra
The goal of the nulling methods we are considering is to mitigate the suppression on the cosmic shear signal due to baryon feedback, shown in Figure 4 for the BNT transformed (blue-dashed lines) and the de-correlated (red-dotted lines) cases. The BNT transformed case is only shown for the auto-correlations and close pairs of tomographic bins as the well separated bins have close to zero correlation by design. The ‘x-cor’ case is shown for the highest three redshift bins as the lowest two redshift bins are used as the lens sample and the third bin is neglected to reduce overlap between lens and source galaxies. The impact for the LUnul method is not shown here as the LUnul method operates in a transformed space (as described in Equation 5), with elements that are no longer indexed by the bin pair indices shown in Figure 4. Instead, we plot the sensitivity to high $k$ effects in Figure 6 for the LUnul method.
@@
Figure 5: Power spectra for each pair of tomographic bins including contribution from different ranges in $k$: the dashed line corresponds to the smallest scales with contributions from $k>0.5h/{\rm Mpc}$, the solid line corresponds to slightly larger scales $(0.1<k<0.5)h/{\rm Mpc}$ and the dotted line shows the full spectra including all $k$. The ‘standard’ case (orange) shows small scales contribute to even the smallest $\ell$-modes, whilst the BNT-transformed spectra (blue) have a more direct mapping between $k$ and $\ell$, especially for the higher redshift bins. The cross-correlation-modified spectra (red, highest three source bins only) exhibit a damping of the contribution from small scales relative to the ‘standard’ cosmic shear spectra; this is most significant in the highest redshift bin which has the largest separation to the lens sample. In the cross-correlation case there are ‘wiggles’ visible in the range $(0.1<k<0.5)h/{\rm Mpc}$ (red solid line) which come from the first lens bin and is due to the presence of the BAO in the clustering signal.
@@ 6 Impact of nulling strategies on cosmological constraints
where $\chi^{ij}$ is the comoving distance at the median of the lensing kernel of the lowest redshift bin in the pair $(ij)$ – for the BNT method we use the transformed kernel. The $\ell$ cut for a given $k_{\rm max}$ is shown for each tomographic bin in Figure 7. There are several alternative ways in which scale-cuts can be implemented, such as a redshift dependent $k$-cut or different $\ell$-cut per tomographic bin pair (desy3/etal:2022). As we do not seek to find optimal scale cuts in this analysis but to simply compare different methods we have only considered one approach, however this choice may impact the results reported here.
@@
where $C_{\ell}^{\rm sys}$ is the change due to baryons. The Fisher matrix (1997ApJ...480...22T) is given by
@@
The parameter bias (2008MNRAS.391..228A) is computed from the Fisher matrix and the bias vector $B_{j}$ as
@@
For all nulling methods, a higher $k_{\rm max}$ leads to a decrease in the error bar but also a larger bias in the parameters, as expected. Unsurprisingly, $S_{8}$, which approximately quantifies the amplitude of the power spectrum, is the most strongly biased parameter (due to unmodelled baryon feedback) in all cases. A bias of less than $1\sigma$ in this parameter is only achieved for the most aggressive scale cut ($k_{\rm max}=0.1\,h/{\rm Mpc}$) in every method apart from the LUnul method; for this method, the bias in $S_{8}$ can be reduced to less than $0.5\sigma$ for a less aggressive scale cut, $k_{\rm max}=0.3\,h/{\rm Mpc}$. The cost of this, however, is that the constraining power is reduced to a similar level as the standard case at $k_{\rm max}=0.1\,h/{\rm Mpc}$. Considering constraints on $S_{8}$ alone, it is therefore difficult to motivate using any proposed nulling method to improve accuracy without the cost of lost precision. This statement only applies to attempts to remove sensitivity to all baryon feedback from the data vector however; if we are willing to model feedback then biases could be reduced and nulling methods may still offer significant benefits to guard against model inaccuracies on small scales. We have focused on the extreme case of ignoring baryon feedback altogether in the model, since we consider it desirable to construct a data vector entirely immune to small-scale astrophysics in order to demonstrate robustness of cosmological parameter constraints from cosmic shear.
We find only small differences between the standard and BNT approaches, particularly in $S_{8}$. Constraints on $w_{0}$ and $w_{a}$ are less biased than the standard case for higher values of $k_{\rm max}$ although the FoM remains comparable across scale-cuts, as shown in the lower panel of Figure 10. This is likely due to there being only six tomographic bins, as previous work (gu/etal:2025) has shown that the BNT method is much more effective when there are more redshift bins, which allows for a cleaner separation of the lens redshifts that contribute to each tomographic shear signal.
@@
For all parameters, we find that that there is a nulling method which can improve parameter constraints in terms of both accuracy and precision. The BNT method is advantageous for $w_{0}-w_{\rm a}$, the cross-correlation approach reduces absolute biases but increases biases relative to uncertainties, and the LUnul method is successful at removing biases but its numerical implementation needs to be tuned appropriately.
@@ 7 Conclusions
LUnul is consistently the most robust method, allowing to push to smaller scales compared to the other methods, as we prioritise reducing parameter bias. Whilst this method consistently achieves a lower FoB in both $S_{8}-\Omega_{\rm m}$ and $w_{0}-w_{\rm a}$, the FoM is always lower than other methods. We do not find a large difference between the standard and BNT constraints, this is likely due to there being only 6 redshift bins in our fiducial setup (a realistic choice for Stage-IV surveys, e.g. 2025arXiv250107559W), where previous works (gu/etal:2025) have forecasted that the BNT method has a bigger impact when considering a large number of well behaved redshift bins. The cross-correlation approach could potentially be set-up more optimally with a different lens selection to maximise how much of the data is used whilst also reducing the amount of overlap between lenses and sources. This approach likely works more effectively for CMB lensing since the lensing kernel peaks at around a redshift of two and thus the lens source separation is greater. Additionally we have assumed linear galaxy bias in this case, which reduces dependence on galaxy bias because of how the estimator is defined. This assumption may not be sufficient given the goal here is to push down to smaller scales.
We have found here that the LUnul method provides consistently less biased constraints and has potential for being applied beyond a cosmic shear only analysis and with additional complexities like intrinsic alignments. The BNT method is likely more suited to the later data sets from Stage-IV surveys that will benefit from a higher number of tomographic bins. Whilst we have shown that the cross-correlation approach is able to reduce some of the suppression due to baryons, the impact is still present in larger scales due to the overlap between source and lens bins, and therefore refinement in lens selection is required. This analysis adds to the evidence that unmodelled baryon feedback leads to biased cosmological constraints and Stage-IV surveys will need to consider strong scale cuts regardless of nulling choice, or rely on accurate models of baryon feedback. These methods do however provide an alternative to the many options now available for modelling the impact of baryons.
@@ Acknowledgements
Appendix A Additional Figures

### v2

@@ 5 Impact of nulling strategies on the cosmic shear power spectra
The goal of the nulling methods we are considering is to mitigate the suppression on the cosmic shear signal due to baryon feedback, shown in Figure 4 for the BNT transformed (blue-dashed lines) and the de-correlated (red-dotted lines) cases. The BNT transformed case is only shown for the auto-correlations and close pairs of tomographic bins as the well separated bins have close to zero correlation by design. The ‘x-cor’ case is shown for the highest three redshift bins as the lowest two redshift bins are used as the lens sample and the third bin is neglected to reduce overlap between lens and source galaxies  [^4] We tested this choice by checking the impact on cosmological parameter constraints when including the third redshift bin shown in 1 in the lens sample and found the constraint on $S_{8}$ to be biased by at least an additional $0.5\sigma$ compared to the two lens bin case.. The impact for the LUnul method is not shown here as the LUnul method operates in a transformed space (as described in Equation 5), with elements that are no longer indexed by the bin pair indices shown in Figure 4. Instead, we plot the sensitivity to high $k$ effects in Figure 6 for the LUnul method.
@@
Figure 5: Ratio of power spectra for each pair of tomographic bins, showing the contribution from a given range in $k$ to the overall power spectrum: the dashed line corresponds to the smallest scales with contributions from $k>0.5h/{\rm Mpc}$, the solid line corresponds to slightly larger scales $(0.1<k<0.5)h/{\rm Mpc}$ and the dotted line shows the full spectra including all $k$. The ‘standard’ case (orange) shows small scales contribute to even the smallest $\ell$-modes, whilst the BNT-transformed spectra (blue) have a more direct mapping between $k$ and $\ell$, especially for the higher redshift bins. The cross-correlation-modified spectra (red, highest three source bins only) exhibit a damping of the contribution from small scales relative to the ‘standard’ cosmic shear spectra; this is most significant in the highest redshift bin which has the largest separation to the lens sample. In the cross-correlation case there are ‘wiggles’ visible in the range $(0.1<k<0.5)h/{\rm Mpc}$ (red solid line) which come from the first lens bin and is due to the presence of the BAO in the clustering signal.
@@
Fiducial LUnul Transformed
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ 6 Impact of nulling strategies on cosmological constraints
where $\chi^{ij}$ is the comoving distance at the median of the lensing kernel of the lowest redshift bin in the pair $(ij)$ – for the BNT method we use the transformed kernel. The $\ell$ cut for a given $k_{\rm max}$ is shown for each tomographic bin in Figure 7. There are several alternative ways in which scale-cuts can be implemented, such as a redshift dependent $k$-cut or different $\ell$-cut per tomographic bin pair (Abbott et al., 2022). As we do not seek to find optimal scale cuts in this analysis but to simply compare different methods we have only considered one approach, however this choice may impact the results reported here.
@@
where $C_{\ell}^{\rm sys}$ is the change due to baryons. The Fisher matrix (Tegmark et al., 1997) is given by
@@
The parameter bias (Amara and Réfrégier, 2008) is computed from the Fisher matrix and the bias vector $B_{j}$ as
@@
For all nulling methods, a higher $k_{\rm max}$ leads to a decrease in the error bar but also a larger bias in the parameters, as expected. Unsurprisingly, $S_{8}$, which approximately quantifies the amplitude of the power spectrum, is the most strongly biased parameter (due to unmodelled baryon feedback) in all cases. A bias of less than $1\sigma$ in this parameter is only achieved for the most aggressive scale cut ($k_{\rm max}=0.1\,h/{\rm Mpc}$) in every method apart from the LUnul method; for this method, the bias in $S_{8}$ can be reduced to less than $0.5\sigma$ for a less aggressive scale cut, $k_{\rm max}=0.3\,h/{\rm Mpc}$. The cost of this, however, is that the constraining power is reduced to a similar level as the standard case at $k_{\rm max}=0.1\,h/{\rm Mpc}$, this is mainly due to this method being a more aggressive loss of information compared to other methods. Considering constraints on $S_{8}$ alone, it is therefore difficult to motivate using any proposed nulling method to improve accuracy without the cost of lost precision. This statement only applies to attempts to remove sensitivity to all baryon feedback from the data vector however; if we are willing to model feedback then biases could be reduced and nulling methods may still offer significant benefits to guard against model inaccuracies on small scales. We have focused on the extreme case of ignoring baryon feedback altogether in the model, since we consider it desirable to construct a data vector entirely immune to small-scale astrophysics in order to demonstrate robustness of cosmological parameter constraints from cosmic shear.
We find only small differences between the standard and BNT approaches, particularly in $S_{8}$. Constraints on $w_{0}$ and $w_{a}$ are less biased than the standard case for higher values of $k_{\rm max}$ although the FoM remains comparable across scale-cuts, as shown in the lower panel of Figure 10. This is likely due to there being only six tomographic bins, as previous work (Gu et al., 2025) has shown that the BNT method is much more effective when there are more redshift bins, which allows for a cleaner separation of the lens redshifts that contribute to each tomographic shear signal.
@@
For all parameters, we find that there is a nulling method which can improve parameter constraints in terms of both accuracy and precision. The BNT method is advantageous for $w_{0}-w_{\rm a}$, the cross-correlation approach reduces absolute biases but increases biases relative to uncertainties, and the LUnul method is successful at removing biases but its numerical implementation needs to be tuned appropriately. Our results are in agreement with previous work considering the impact of baryonic feedback on cosmic shear analyses. In all cases we recover the now well established trade-off between reducing bias in cosmological constraints versus the achievable statistical precision. We find that aggressive mitigation strategies (whether via scale cuts or nulling methods) can be effective at removing the impact of small-scale systematics like baryon feedback, but at the cost of significantly reducing constraining power. Additionally we see that the $S_{8}$ parameter is the most sensitive to baryonic effects, consistent with previous forecasts and survey analyses.
@@ 7 Conclusions
LUnul is consistently the most robust method, allowing to push to smaller scales compared to the other methods, as we prioritise reducing parameter bias. Whilst this method consistently achieves a lower FoB in both $S_{8}-\Omega_{\rm m}$ and $w_{0}-w_{\rm a}$, the FoM is always lower than other methods. We do not find a large difference between the standard and BNT constraints, this is likely due to there being only 6 redshift bins in our fiducial setup (a realistic choice for Stage-IV surveys, e.g. Wong et al. 2025), where previous works (Gu et al., 2025) have forecasted that the BNT method has a bigger impact when considering a large number of well behaved redshift bins. The cross-correlation approach could potentially be set-up more optimally with a different lens selection to maximise how much of the data is used whilst also reducing the amount of overlap between lenses and sources. This approach likely works more effectively for CMB lensing since the lensing kernel peaks at around a redshift of two and thus the lens source separation is greater. Additionally we have assumed linear galaxy bias in this case, which reduces dependence on galaxy bias because of how the estimator is defined. This assumption may not be sufficient given the goal here is to push down to smaller scales.
We have found here that the LUnul method provides consistently less biased constraints and has potential for being applied beyond a cosmic shear only analysis and with additional complexities like intrinsic alignments. The BNT method is likely more suited to the later data sets from Stage-IV surveys that will benefit from a higher number of tomographic bins. Whilst we have shown that the cross-correlation approach is able to reduce some of the suppression due to baryons, the impact is still present in larger scales due to the overlap between source and lens bins, and therefore refinement in lens selection is required. This analysis adds to the evidence that unmodelled baryon feedback leads to biased cosmological constraints and Stage-IV surveys will need to consider strong scale cuts regardless of nulling choice, or rely on accurate models of baryon feedback. These methods do however provide an alternative to the many options now available for modelling the impact of baryons. The nulling approaches consider in this work may be also useful for reducing the impact of other systematics such as intrinsic alignment and photometric redshift calibration uncertainties.
@@ Acknowledgements
Appendix A Additional Figures and Tables
@@
Method Assumptions Advantages Limitations
BNT Transform shear field to localise lensing kernel in redshift (improves $k\ell$ mapping) Accurate redshift distributions. Cleaner separation of scales. Limited improvement with few bins.
LUnul Apply linear transformation (e.g. LU decomposition) to power spectra to cancel high-$k$ contributions Numerical implementation choices for discretised Limber integral. Generalisable beyond lensing kernels. Harder to extend beyond 2-point/Limber regime.
Cross-correlation Use shear–density cross-correlations to remove low-redshift contributions Linear galaxy bias. Alternative to $3\times 2$pt analyses. Dependent on lens selection and requires additional data and modelling.
Table 1: Summary comparison of nulling methods for mitigating small-scale contributions in cosmic shear analyses.